In a multiple regression analysis, two independent variables are considered, and the sample size is 26. The regression coefficients and the standard errors are as follows.

b1 = 2.312 Sb1 = 0.81
b2 = −1.950 Sb2 = 0.77

Conduct a test of hypothesis to determine whether either independent variable has a coefficient equal to zero. Would you consider deleting either variable from the regression equation? Use the 0.05 significance level. (Negative amounts should be indicated by a minus sign. Round your answers to 3 decimal places.)

H0: β1 = 0 H0: β2 = 0
H1: β1 ≠ 0 H1: β2 ≠ 0
H0 is rejected if t < −2.069 or t > 2.069
t=___
t=___
variable can be deleted.
  • The first and second
  • Neither the first nor second
  • The second
Answer #1

for first variable : t =(coefficient/standard error) =(2.312/0.81)= 2.854

for second variable : t =(coefficient/standard error) =(-1.950/0.77)= -2.532

since both test statistic falls in rejection region

Neither the first nor second